OpenStudy (anonymous):

Find the average value of y = x3 over the interval [0, 2]. Answer: 2??

OpenStudy (anonymous):

x^3

OpenStudy (anonymous):

integral x^3dx divided by integral xdx when x runs from 0 to 2

OpenStudy (anonymous):

that is [x^4/4] when x runs from 0 to 2 divided by [x^2/2] when x runs from 0 to 2

OpenStudy (anonymous):

= [2^4/4 -0]/[2^2/2-0] = 4-2 =2